OpenStudy (anonymous):

factor 4x^4-2x^3y-3xy^3-9y^4

OpenStudy (victoriasushchik):

do you mean this: http://prntscr.com/aba2ve

OpenStudy (anonymous):

yes

OpenStudy (kittiwitti1):

\[4x^{4}-2x^{3}y-3xy^{3}-9y^{4}\]rearranging it\[4x^{4}-9y^{4}-2x^{3}y-3xy^{3}\]group method:\[\left(2x^{2}-3y^{2}\right)^{2}-xy\left(2x^{2}-3y^{2}\right)\]factor out 2x^2-3y^2 from both:\[\left(2x^{2}-3y^{2}\right)\left[\left(2x^{2}-3y^{2}\right)-xy\right]\]

OpenStudy (kittiwitti1):

Still confused?

OpenStudy (kittiwitti1):

Hey, if you're not fully understanding this I can explain to you. That's fine with me ^_^